CSePub Home Book Store Instructors Independent Authors Ab You have saved S23695 00for athree-year vacation to Germany. You will keep your money invested in an account paying 804% APR with monthly compounding while you live in Munich You will make your withdrawals at the beginning of the month What is the balance on the account after the first month? (This is the balance prior to the withdrawal for the second month) Answer Format: Currency: Round to: 2 decimal places Enter Answer Here Submit Answer Prev Problem 1Al Problems Ned Contact Us Privacy Puy T
Answer #1

Answer:

Saving amount = PV = $23,695

Monthly Interest = 8.04%/12 = 0.67%

Number of monthly withdrawals = 3 *12 = 36

Withdrawals are made at beginning of the month.

First let us calculate monthly withdrawal amount.

To get monthly withdrawal, we will use PMT function of excel:

PMT (rate, nper, pv, fv, type)

PMT (0.67%, 36, -23695, 0, 1)

=$738.01

Starting balance = $23,695

Withdrawal Month 1 = $738.01

Interest of month 1 = ($23,695 - $738.01) * 0.67% = $153.81

Hence,

Balance in account after the first month = $23,695 - $738.01 +  $153.81 = $23110.80

Balance in account after the first month = $23,110.80

   $23,110.80